andy has eighteen coins made up of quarters and half dollars, and their total value is $7.50. How many quarters does he have?
Answer by alicealc(293)

number of quarters coins = q
number of half-dollars coins = h

q + h = 18

1/4 q + 1/2 h = 7.5
multiply all by 4:
q + 2h = 30

use the elimination procedure:
q + h = 18
q + 2h = 30
------------ (-)
-h = -12
h = 12

q + 12 = 18
q = 18 - 12 = 6

So he has 6 quarters